OpenStudy (anonymous):

what is the standard form of the equation of the circle 3x^2+3y^2-12y-0?

OpenStudy (anonymous):

i assume that is \(=0\) not \(-0\)

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

divide by 3 first and start with \[x^2+y^2-4y=0\] then complete the square for the \(y\) terms do you know how to do that?

OpenStudy (anonymous):

No I do not

OpenStudy (anonymous):

what is half of \(4\) ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok then we start with \[x^2+(y-2)^2=?\] now what is \(2^2\) ?

OpenStudy (anonymous):

I thought it was a -4?

OpenStudy (anonymous):

complete the square for y... (-4/2)^2

OpenStudy (anonymous):

minus does not matter when you square it

OpenStudy (anonymous):

half of \(-4\) is \(-2\) and then \(2^2=4\) so you get \[x^2+(y-2)^2=4\]